OpenStudy (anonymous):

A truck is moving 14 m/s to the south. If it takes the truck 20 seconds to reduce its velocity to 10 m/s south, what is its acceleration?

OpenStudy (abhisar):

In the question, we are given with, Initial velocity = 14m/s Final velocity = 10m/s Time taken = 20 s Now, by using the equation of motion, v=u+at 10 = 14 + a \(\times\) 20 \(\Rightarrow\) -4 = 20a a= ?

OpenStudy (shamim):

U can also use a=(v-u)/t

OpenStudy (shamim):

Both equations r identical